OpenStudy (anonymous):

hellllpppp please 5x^2+1)(6x^2+7)

OpenStudy (anonymous):

do you know how to expand (a+b)(c+d) ?

OpenStudy (anonymous):

no i need help

OpenStudy (anonymous):

do you know that a(c+d) = ac+ad?

OpenStudy (anonymous):

well to expand you multiply each term in the first set of parenthesis by each term in the second set of parenthesis so (a+b)(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

i did that i stil go the wrong answer

OpenStudy (amistre64):

What answer do you get? perhaps the material you are using is in error

OpenStudy (anonymous):

yeah what are you getting?